OpenStudy (anonymous):

for f(x)=3x+1 and g(x)=x^2-6, find (f o g)(4)

OpenStudy (anonymous):

31

OpenStudy (amistre64):

determine the value of g(x) when x=4; then apply that value from g into f(x)

OpenStudy (amistre64):

(f o g) simply means f[g(x)] where g(x) is the value of f's stuff

OpenStudy (amistre64):

they prolly should use different variables in them to lessen any confusions perhaps

OpenStudy (bahrom7893):

they should and I never understand (f o g), they should've just said: f(g(x))

OpenStudy (amistre64):

i blame leibnez :)

OpenStudy (anonymous):

would you all agree with the answer of 31 from jkip

OpenStudy (amistre64):

lets see: (4)^2 - 6 = ?

OpenStudy (anonymous):

10...?

OpenStudy (amistre64):

10 is correct :) 3(10) + 1 = ?

OpenStudy (anonymous):

g(4)=4^2-6=10 then F(10)=3*10+1=31

OpenStudy (anonymous):

thankyou both